OpenStudy (anonymous):

Three out of 200 machine parts were found to be defective. About how many parts would be expected to be defective out of 2500 parts

OpenStudy (anonymous):

what percentage of 200 is three?

OpenStudy (anonymous):

3/200

OpenStudy (calculusxy):

I think that a proportion can work. \[\frac{ 3 }{ 200 } = \frac{ x }{ 2500 }\]

OpenStudy (anonymous):

Ok so you would take that percentage and multiply it by 2500 and get 37.5 round up to 38

OpenStudy (calculusxy):

We have the sample proportion as: \[\frac{ defective }{ total }\]

OpenStudy (calculusxy):

Now what you just need to do is solve for x.

OpenStudy (anonymous):

@Directrix

Directrix (directrix):